OpenStudy (rachie19):

2x/x-2 -2/x+1=4 solve for x help please =(

OpenStudy (xguardians):

Please write the equation with proper parenthesis.

rebeccaxhawaii (rebeccaxhawaii):

click "equation" and write it out.

OpenStudy (rachie19):

\[\frac{ 2x }{ x-2 }-\frac{ 2 }{ x+1 }=4\]

OpenStudy (rachie19):

@rebeccaxhawaii

OpenStudy (rachie19):

@xGuardians

OpenStudy (anthonyym):

You want to have a common denominator first. A common denominator would be (x-2)(x+1)

OpenStudy (anthonyym):

So multiply the first fraction by (x+1)/(x+1) and the second fraction by (x-2)/(x-2).